OpenStudy (anonymous):

How do i express this a ratio of integers? 6.254545454545454.........

OpenStudy (mertsj):

1. Write the equation: n=6.254545... 2. Multiply that equation by 100 because there are two digits that repeat and 100 has two zeros. 3. Write the new equation ABOVE the Step 1 equation. 4. Subtract the bottom equation from the top. 5. Solve the resulting equation for n

OpenStudy (anonymous):

n x 100 = 625.45454....

OpenStudy (anonymous):

625.4545454... 6.254545454... - ________________

OpenStudy (mertsj):

100n=625.454545... n= 6.254545...

OpenStudy (anonymous):

subtract what you have written above?

OpenStudy (mertsj):

yes. And notice that you end up with a string of zeros.

OpenStudy (mertsj):

99n=619.20000000000....

OpenStudy (anonymous):

99n= 619.20000000...

OpenStudy (anonymous):

Yep

OpenStudy (mertsj):

Divide both sides by 99 and drop the zeros

OpenStudy (mertsj):

\[n=\frac{619.2}{99}\]

OpenStudy (anonymous):

6.245454545...

OpenStudy (mertsj):

Now multiply that fraction by 10 over 10 to get rid of the decimal and then reduce the fraction.

OpenStudy (anonymous):

that's what we did before too... correct?

OpenStudy (mertsj):

\[n=\frac{619.2}{99}\times \frac{10}{10}=\frac{6192}{990}\]

OpenStudy (anonymous):

sorry...yes

OpenStudy (anonymous):

the answer in my solution manual is 6.2 +54/10^3 + 54/ 10^5 sorry I should have mentioned that earlier....so how do we get there

OpenStudy (mertsj):

Well I guess I don't understand because the directions say to write the number as a ratio of integers. The answer you gave is NOT a ratio let alone a ratio of 2 integers.

OpenStudy (anonymous):

Is is because the repeating digits are in the thousands position and so on?

OpenStudy (anonymous):

that's what my book says too....so is the solution manual off?

OpenStudy (mertsj):

Yes

OpenStudy (anonymous):

54/(10^3) is not a ratio of integers?

OpenStudy (mertsj):

Yes it is but that is only a small part of the solution you posted.

OpenStudy (mertsj):

It's like saying that 1/2 +3/4 + 1/7 is the ratio of two integers and it clearly isn't.

OpenStudy (anonymous):

I'm sooo sorry....

OpenStudy (anonymous):

here is the complete solution ...I'M SOOO SILLY

OpenStudy (anonymous):

344/55

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

That is the answer.

OpenStudy (mertsj):

\[\frac{6192}{990}=\frac{344}{55}\]

OpenStudy (anonymous):

6.2545454...=6.2+54/10^3 +....= (6.2 + 54/10^3)/(1-(1/10^2))= 62/10 +54/990= 6192/990= 344/55

OpenStudy (anonymous):

THANKS SOO MUCH....makes perfect sense now....

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

cool thanks. soo.... 1. Write the equation: n=6.254545... 2. Multiply that equation by 100 because there are two digits that repeat and 100 has two zeros. 3. Write the new equation ABOVE the Step 1 equation. 4. Subtract the bottom equation from the top. 5. Solve the resulting equation for n use these instructions whenever i need to change and expression into a ratio of integers